log 3log 9+log 729+3log √3³ find the value please send me answer

Grade:12

1 Answers

Sher Mohammad IIT Delhi
10 years ago
3log3*3=6log3 729=3^6
log729=6log3
log v3³=(3/2)log3=1.5log3

3log 9+log 729+3log v3³=6log3+6log3+4.5log3=16.5log3
if base is 3, than answer is 16.5



Sher Mohammad
Name,
askIITians Faculty
B.Tech.
IIT Delhi